Trevor and Marissa together have 26 T-shirts to sell. If Marissa has 6 fewer T-shirts than Trevor, find how many T-shirts Trevor has. Use x as your variable.

Answer by ikleyn(52794)

.

    x + y = 26     (1)  (x = # of T-shirts Trevor has to sell)

    x - y =  6     (2)


Add the equations and get


    2x     = 26 + 6 = 32;   x = 32/6 = 16.


Then from equation (1)  y = 26-16 = 10.


ANSWER.  Trevor has 16 T-shirts;  Marissda has  10.

Solved.

------------

See a TWIN problem solved under this link

https://www.algebra.com/algebra/homework/word/finance/Money_Word_Problems.faq.question.1172822.html

https://www.algebra.com/algebra/homework/word/finance/Money_Word_Problems.faq.question.1172822.html

one hour ago.


Also,  see a million similar solved problems in the lesson
    - Word problems that lead to a simple system of two equations in two unknowns
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Systems of two linear equations in two unknowns".

Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.